A circle has an area of 841pi cm squared. What is the circumference?
Answer by mananth(16946)

Area = 841 pi
841pi = pi*r^2
841=r^2
r= sqrt(841)=29
Circumference = 2*pi*r
Circumference = 2*pi*29
=182.2 cm